# TBRD: TESLA Authenticated UAS Broadcast Remote ID

### v1

@@ V Security and Performance Analysis
UAS compromise N/A ✗ long term private key compromise ✓key compromise for current mission only
@@
Real-time Verification N/A Difficult
@@
V-A Spoofing-Replay-Timing Attack Resilience
We now evaluate the types of attacks that can be launched during an active UAS mission and the extent to which an attacker could spoof the existence of a drone that doesn’t actually exist, replay previously captured TBRD messages, or attack the system by targeting the time synchronization parameters of the TESLA protocol. In scenarios where an attacker launches attacks against a legitimate operator, we assume the attacker knows the operator ID, UAS ID, and has sufficient advanced notice of the impending mission (perhaps even an exact start time). The attacker is located in close proximity to the UAS mission, can effectively receive TBRD messages, and even broadcast with sufficient power to overshadow or jam legitimate transmissions. The TESLA protocol is a key feature of TBRD that protects against attacks during active missions. The experimental results presented in Section IV demonstrate the security guarantees provided by the TBRD protocol. However, two potential weaknesses that can be targeted to affect the security guarantees of TESLA are the key disclosure period and time synchronization.
The key disclosure period determines the time during which a key remains secret. Once a key is disclosed, the observer can validate previously received messages. The time between message broadcast and key disclosure allows an attacker to inject fake or manipulated data. During this interval, care must be taken not to trust received data implicitly. The TESLA parameters can be tuned such that this interval is minimized at the cost of a larger number of interval keys, which may impact the total flight time of a TBRD-protected mission. For example, suppose an observer intends to use TBRD messages for real-time collision avoidance. In that case, they may wish to minimize the opportunity for an adversary to send fake location data. Conversely, the TESLA parameters may be chosen so that the system design can withstand the impact of an attacker spoofing or manipulating information during a known interval. TBRD can be customized to allow operators to weigh the most relevant risks.
@@ V-A Spoofing-Replay-Timing Attack Resilience
Figure 9: Fake Message Injection Within Key Interval. The attacker can inject a fake message during the current key interval, and forgery is detected by the observer in the subsequent interval.
@@
TESLA requires loose time synchronization between transmitter and receiver to maintain security guarantees [9]. When a receiver wishes to verify the authenticity of a received message, the primary goal is to ensure that the key used to calculate the HMAC of the payload was still a secret at the time of HMAC generation. In other words, the receiver must not trust a message that uses a key that has already been disclosed. Thus, time synchronization between the sender and receiver must be maintained to ensure all parties are in lock-step with when the key has been used and when a new key is disclosed. Our system assumes GPS time to synchronize the time between the transmitter and the receivers. Spoofing or jamming GPS as a mechanism to attack the TESLA protocol is beyond the scope of this work. However, we acknowledge that even with all parties using GPS synchronized clocks, there will still be some clock skew. One relevant work investigated the impacts of GPS clock synchronization on VANETs implementing TESLA authentication. It was determined that even with a maximum clock skew of 1.5 ms, there is potential for an attack if additional protections are not implemented  [28]. Beyond the clock skew introduced by GPS, special care must be taken to account for wireless broadcast protocols that may delay the transmitter’s ability to broadcast messages within the intended interval. As we discussed in Section IV, we propose using Wi-Fi beacons within a permitted transmission interval related to the interval period. This ensures that the media access control mechanisms imposed by the 802.11 protocol will not negatively affect the security guarantees of our proposed system. As each wireless broadcast mechanism employs a different media access control mechanism (e.g., Bluetooth, cellular), the parameters of TBRD must be modified to account for varying applications.
@@
V-B Computation and Communication Overhead
@@
Computation Analysis: To validate claims that the TBRD protocol provides reduced power consumption when compared to authentication through digital signatures, we developed an experimental method to examine the power necessary for signature generation as a measure of computational efficiency. As many UAS must operate under strict battery power limitations, power consumption is a key concern. For our analysis, we must ensure that we compare algorithms that provide a similar level of security. While the security strength of cryptographic algorithms depends on many factors, we can rely on well-established algorithms with clear standards to ensure a systematic approach. One such standard is NIST Special Publication 800-57 Part 1 [44]. This standard compares various cryptographic algorithms and assigns a security strength (measured in bits) as a function of the algorithm type and key size. Our implementation of the TBRD protocol utilizes a SHA-256 HMAC to achieve authentication, providing 256 bits of security strength. To provide an equivalent level of security (256 bits), an elliptic-curve cryptography (ECC) key size of at least 512 bits is required [44].
@@
We chose to conduct our tests of cryptographic power consumption using the ESP32-S3-DevKitM-1 [45]. This platform was chosen due to its ease of use, support for rapid prototyping and testing, as well as the prevalence of ESP32-based Remote ID transponders already in everyday use. As mentioned previously, ArduPilot has spearheaded the open-source implementation of a Remote ID-compliant transmitter. The list of supported hardware includes seven boards, including the ESP32-S3 dev board, and several other ESP32-based solutions [32, 46, 47]. The boards offer a compact, lightweight form factor for UAS operators to retrofit existing UAS platforms with Remote ID compliant transmitters.
@@
To isolate the effects of the message authentication process, we programmed our ESP32-based microcontrollers to calculate a SHA256 HMAC of a 100-byte payload once every second using the WolfSSL embedded cryptography library [48]. The 100-byte payload mirrors the dynamic flight data, used to generate the HMAC used by the TBRD protocol, and the signature rate complies with the ASTM standard broadcast rate for Remote ID. Using a Nordic Semiconductor Power Profiler II device, we were then able to measure both the computation time and the amount of power used during the HMAC calculations [49]. Similarly, we can analyze the calculation of an ECC digital signature for a 100-byte payload with a 512-bit key size. However, the ESP32-S3 was unable to compute the digital signature at a rate of once per second. While other microcontrollers may be able to compute the digital signature at a faster rate, the configuration we tested would not be able to comply with the minimum transmission rate specified by the Open Drone ID standard [50]. Figure 10 depicts the power consumption used by the microcontroller during computational intervals. On average, the SHA256 HMAC took 10 ms while the ECC digital signature took 1.2 seconds. Additionally, it is worth noting that the SHA256 HMAC generates a 32-byte hash output, whereas the ECC algorithm produces a 139-byte signature. The average power required for a SHA256 HMAC is 129 mW, while the average power needed for an ECC digital signature is 216 mW.
@@
Figure 10: Power consumption of an ESP32-S3 microcontroller computing the ECC digital signature of a 100-byte payload compared to the same device computing a SHA256 HMAC.
@@
Communication Overhead: There are several constraints we must assess to evaluate the performance of the TBRD protocol. Since we have chosen to implement TBRD within the framework of the ASTM Open Drone ID standard, we must ensure that the additional data for authentication conforms to the relevant message size constraints. The standard limits the total length of all authentication/signature data to 255 bytes. The TBRD authentication data is 68 bytes in length when using a SHA-256 HMAC for authentication. It is essential to note that the authentication data for TBRD may be compromised with trade-offs to system security. The SHA256 HMAC provided 256 bits of security, while the TESLA protocol by Galileo operates with a truncated HMAC of only 40 bits [10]. Again, if we compare the use of an equally secure digital signature scheme like ECC, the signature size would double the necessary length of the authentication data. When operating in a wireless environment with a high potential for packet loss, it is crucial to minimize the overhead required for message authentication. TBRD excels in this regard as the same message size can be authenticated with a smaller amount of data compared to a digital signature approach.
VI Discussion
@@
TBRD Parameter Configuration: The TBRD protocol allows for many adjustments to system parameters to enable trade-offs in security and performance. The parameters include the decision in the delay disclosure period, the interval period, and the adjustable HMAC size. The TESLA protocol introduces a controlled delay $d$ between the time a message is sent and when the authentication key is disclosed. This delay must be carefully balanced against real-time requirements. In latency-sensitive applications (e.g., drone command and control), a long delay may degrade responsiveness, while a shorter delay could reduce the time available for key distribution and verification. The interval duration $T_{int}$ must also be tuned to match message transmission frequency and expected network conditions. For example, high message frequency may require shorter $T_{int}$ to minimize overlap and reduce buffer size requirements. Conversely, longer intervals may reduce overhead but increase exposure to packet loss or adversarial manipulation. Careful parameter selection is crucial for optimizing both security and operational performance. Lastly, the HMAC output may be truncated to address constraints resulting from even more lossy wireless communication channels.
@@
Mobile compromise: In TBRD, the key commitment of the TESLA-based broadcast authentication depends on the mobile device to serve as a trusted execution platform for key disclosure and signature operations. However, the integrity of this process hinges on the security guarantees of the TEE itself.. Current TEEs (e.g., ARM TrustZone or Apple Secure Enclave) offer limited visibility and control to developers, and vendor-specific APIs or OS policies often constrain their use. Moreover, if the mobile device is compromised (e.g., rooted or jailbroken), the adversary may gain access to pre-disclosure keys. This highlights a significant limitation: the security model assumes the mobile TEE is trustworthy, but practical deployments must account for the possibility of device-level compromise.
@@
Standards integration: Integrating TESLA into the ASTM Remote ID framework presents both opportunities and challenges. TESLA offers a lightweight, broadcast-friendly authentication primitive that complements the ASTM requirement for message authenticity without necessitating public key infrastructure on resource-constrained drones. However, compliance with current ASTM standards (e.g., F3411) requires strict adherence to message construction and transmission. Furthermore, compatibility with transmission mechanisms (Bluetooth, Wi-Fi) adds additional constraints.
VII Related Work
@@
Several works have focused on the privacy and security issues associated with compliance with the Remote ID rule. In  [51], Tedeschi et al. provide perhaps the most comprehensive overview of privacy and security issues facing Remote ID compliance by UAS operators and manufacturers. Their work identifies several potential threats and solutions, considering the security, privacy, and energy trade-offs associated with these solutions. While a wide variety of technical approaches were considered, their work did not address or mention the use of the TESLA algorithm.
@@
Of the numerous works that propose a technical solution or protocol to enhance Remote ID security, the most closely related is CAPSID [52]. In their work, Li et al. propose the CAPSID protocol, which builds upon the proposed use of session IDs. At its core, CAPSID utilizes certificates to authenticate broadcast messages. Our solution leverages the TESLA protocol to achieve authentication in a more computationally efficient manner. Additionally, they do not provide a practical solution to UAS capture. TBRD ensures that capture of a UAS does not give an adversary access to long-term key material. Furthermore, other closely related works attempt to provide solutions that enhance operator privacy. Privacy of operator location was not a focus of our work [52, 53, 54].
@@
The application of TESLA to solve current and future issues facing broadcast authentication is not unique. TESLA is in use by the Galileo global navigation satellite sytem [55]. TESLA was proposed as a solution for providing authentication to Automated Dependent Surveillance-Broadcast (ADS-B) in [20], which is widely used in the aviation industry. For the maritime sector, it has been proposed to enhance authentication of Automated Identification Systems (AIS)  [19] as well as its replacement, VHF Data Exchange (VDES)  [21]. Our system has many similarities with AuthAUS. One key difference is that AuthAIS utilizes a Trusted Third Party Maritime Authority to assign and distribute root keys to vessels that wish to use TESLA-authenticated AIS. They also propose a probabilistic security framework that reduces message overhead. Using TESLA to authenticate messages in these industries has not yet caught on with the mainstream. However, the explosive growth and unique constraints of UAS operations require a novel approach. Our work demonstrates that using the TESLA algorithm combined with trusted computing platforms can overcome previously assumed entry barriers.
VIII Conclusion
@@
Appendix A Usage
There are four main components to the TBRD proof of concept implementation:
  1. 1.
    TBRD Mobile Application
  2. 2.
    USS Server
  3. 3.
    TBRD Transmitter
  4. 4.
    TBRD Receiver
A-A TBRD Mobile Application
The mobile application is an Android app that uses the Android Keystore system, which leverages the mobile device’s Trusted Execution Environment (TEE) to generate the TBRD authentication keys.
Before flying the UAS, the pilot will use the mobile app to enter the mission information that is used to generate the TBRD keys. The pilot will enter the mission start time, mission end time, UAS ID, and Operator ID, and click on the ‘Generate Keys’ button. The app creates a random key in the TEE, which is used to generate all the subsequent keys in the keychain. The initial seed key never leaves the TEE. The system is designed such that once all the mission keys are generated, the app will send the keychain over to the UAS to use during the mission. However, in our proof of concept implementation, the application simply generates a txt file that can be manually transferred to the TBRD transmitter file system.
Figure 11: TBRD Mobile Application
A-B USS Server
Before running the TBRD transmitter, you must start the USS server. This server accepts and processes incoming TBRD messages. The command line options are outlined in Table II.
TABLE II: USS Command-line Options
Argument Type Description Default
–ip String IP address to bind to 0.0.0.0
–port Int Port to bind to 5555
A-C Wireless Network Interface Configuration
Before running either the TBRD transmitter or TBRD receiver, you must first place your wireless interface in monitor mode and configure transmit/receive for a specific wireless channel. If you are using airmon-ng, an example sequence of commands is:
sudo airmon-ng check
sudo airmon-ng check kill
sudo airmon-ng start \$interface
sudo iw dev \$interfacemon set channel 6
sudo ip link set \$interfacemon down
sudo ip link set \$interfacemon up
Figure 12: USS Test
A-D TBRD Transmitter
The TBRD transmitter can broadcast a set of static data (helpful for testing/debugging) or ingest simulated flight data from ArduPilot’s software-in-the-loop (SITL) simulation. See the SITL Simulation Tutorial here for more information on how to set up your simulation before running the TBRD transmitter. The TBRD transmitter application accepts several command-line arguments for configuration as outlined in Table III:
TABLE III: TBRD Tx Command-Line Options
Argument Short Description Default
–static -s Use static data (disabled)
–verbose -v Enable verbose output (disabled)
–udp -u Broadcast TBRD over UDP (disabled)
–config -c Path to configuration file tx_config.ini
Figure 13: TBRD Transmitter Test
A-E TBRD Receiver
The TBRD Receiver listens for TBRD messages over Wi-Fi or UDP, extracts the TBRD authentication payload, and verifies the authenticity of each payload. Command-line options are outlined in Table IV
TABLE IV: TBRD Rx Command-Line Options
Argument Short Description Default
–verbose -v Enable verbose output (disabled)
–udp -u Sniff UDP instead of Wi-Fi (disabled)
–config -c Path to configuration file rx_config.ini
Figure 14: TBRD Receiver Test

### v2

@@ 5.2. Spoofing-Replay-Timing Attack Resilience
UAS compromise N/A ✗ long term private key compromise ✓key compromise only for current mission
@@
Real-time Verification N/A Difficult Easy
@@
As the swarm simulation demonstrated, TBRD’s USS verification mechanism is fully resilient against mission-level spoofing attacks. An adversary broadcasting messages without a valid, USS-registered key commitment will be immediately identified. In this section, we analyze more subtle attacks that target the intrinsic properties of the TESLA protocol itself and its impact on TBRD security guarantees: (1) the temporary window of uncertainty created by delayed key disclosure, and (2) attacks on the system’s time synchronization. The TESLA protocol is a key feature of TBRD that protects against attacks during active missions. The experimental results presented in this section demonstrate the security guarantees provided by the TBRD protocol. However, two potential weaknesses that can be targeted to affect the security guarantees of TESLA are the key disclosure period and time synchronization.
The primary trade-off of TESLA is the key disclosure period, which creates a temporary window of uncertainty. An attacker can inject a fake message during the current key interval. The observer receives this message but cannot validate or reject it until the next interval when the corresponding key is disclosed. While the attack is detected (and the fake message is ultimately discarded) upon key disclosure, the system is blind to the forgery for the duration of that interval. Care must be taken not to trust received data implicitly during this window. For highly sensitive, real-time applications such as collision avoidance, a 1-second delay in verification is a critical operational consideration that system designers must account for and configure according to their specific use case.
@@
Figure 8. Fake Message Injection Within Key Interval. The attacker can inject a fake message during the current key interval, and forgery is detected by the observer in the subsequent interval.: ¡short description¿
@@
TESLA requires loose time synchronization between transmitter and receiver to maintain security guarantees (perrig_tesla_2002). When a receiver wishes to verify the authenticity of a received message, the primary goal is to ensure that the key used to calculate the HMAC of the payload was still a secret at the time of HMAC generation. In other words, the receiver must not trust a message that uses a key that has already been disclosed. Thus, time synchronization between the sender and receiver must be maintained to ensure all parties are in lock-step with when the key has been used and when a new key is disclosed. Our system assumes GPS time to synchronize the time between the transmitter and the receivers. Spoofing or jamming GPS as a mechanism to attack the TESLA protocol is beyond the scope of this work. However, we acknowledge that even with all parties using GPS synchronized clocks, there will still be some clock skew. One relevant work investigated the impacts of GPS clock synchronization on VANETs implementing TESLA authentication. It was determined that even with a maximum clock skew of 1.5 ms, there is potential for an attack if additional protections are not implemented  (jahanian_analysis_2015). Beyond the clock skew introduced by GPS, special care must be taken to account for wireless broadcast protocols that may delay the transmitter’s ability to broadcast messages within the intended interval. As we discussed in Section 4, we propose using Wi-Fi beacons within a permitted transmission interval related to the interval period. This ensures that the media access control mechanisms imposed by the 802.11 protocol will not negatively affect the security guarantees of our proposed system. As each wireless broadcast mechanism employs a different media access control mechanism (e.g., Bluetooth, cellular), the parameters of TBRD must be modified to account for varying applications.
Replay Attacks- True replay attacks (i.e., re-broadcasting an old, valid message) are prevented by two mechanisms. First, the interval counter $i$, included in every authentication payload prevents the re-use of a message within the same mission . Second, the USS mission time check ensures that an attacker cannot replay a captured broadcast from a previous, completed mission .
@@
5.3. Computation and Communication Overhead
@@
Computation Analysis: To validate claims that the TBRD protocol provides reduced power consumption when compared to authentication through digital signatures, we developed an experimental method to examine the power necessary for signature generation as a measure of computational efficiency. As many UAS must operate under strict battery power limitations, power consumption is a key concern. For our analysis, we must ensure that we compare algorithms that provide a similar level of security. While the security strength of cryptographic algorithms depends on many factors, we can rely on well-established algorithms with clear standards to ensure a systematic approach. One such standard is NIST Special Publication 800-57 Part 1 (barker_recommendation_2020). This standard compares various cryptographic algorithms and assigns a security strength (measured in bits) as a function of the algorithm type and key size. Our implementation of the TBRD protocol utilizes a SHA-256 HMAC to achieve authentication, providing 256 bits of security strength. To provide an equivalent level of security (256 bits), an elliptic-curve cryptography (ECC) key size of at least 512 bits is required (barker_recommendation_2020).
@@
We chose to conduct our tests of cryptographic power consumption using the ESP32-S3-DevKitM-1 (espressif_esp32-s3-devkitm-1_2025). This platform was chosen due to its ease of use, support for rapid prototyping and testing, as well as the prevalence of ESP32-based Remote ID transponders already in everyday use. As mentioned previously, ArduPilot has spearheaded the open-source implementation of a Remote ID-compliant transmitter. The list of supported hardware includes seven boards, including the ESP32-S3 dev board, and several other ESP32-based solutions (ardupilot_ardupilot_2025; holybro_remote_2025; blue_mark_dronebeacon_2025). The boards offer a compact, lightweight form factor for UAS operators to retrofit existing UAS platforms with Remote ID compliant transmitters.
@@
To isolate the effects of the message authentication process, we programmed our ESP32-based microcontrollers to calculate a SHA256 HMAC of a 100-byte payload once every second using the WolfSSL embedded cryptography library (wolfssl_wolfssl_2017). The 100-byte payload mirrors the dynamic flight data, used to generate the HMAC used by the TBRD protocol, and the signature rate complies with the ASTM standard broadcast rate for Remote ID. Using a Nordic Semiconductor Power Profiler II device, we were then able to measure both the computation time and the amount of power used during the HMAC calculations (nordic_semiconductors_power_2025). Similarly, we can analyze the calculation of an ECC digital signature for a 100-byte payload with a 512-bit key size. However, the ESP32-S3 was unable to compute the digital signature at a rate of once per second. While other microcontrollers may be able to compute the digital signature at a faster rate, the configuration we tested would not be able to comply with the minimum transmission rate specified by the Open Drone ID standard (segger_emsecure-ecdsa_2025). Figure 9 depicts the power consumption used by the microcontroller during computational intervals. On average, the SHA256 HMAC took 10 ms while the ECC digital signature took 1.2 seconds. Additionally, it is worth noting that the SHA256 HMAC generates a 32-byte hash output, whereas the ECC algorithm produces a 139-byte signature. The average power required for a SHA256 HMAC is 129 mW, while the average power needed for an ECC digital signature is 216 mW.
@@
Figure 9. Power consumption of an ESP32-S3 microcontroller computing the ECC digital signature of a 100-byte payload compared to the same device computing a SHA256 HMAC.: ¡short description¿
@@
Communication Overhead: There are several constraints we must assess to evaluate the performance of the TBRD protocol. Since we have chosen to implement TBRD within the framework of the ASTM Open Drone ID standard, we must ensure that the additional data for authentication conforms to the relevant message size constraints. The standard limits the total length of all authentication/signature data to 255 bytes. The TBRD authentication data is 68 bytes in length when using a SHA-256 HMAC for authentication. It is essential to note that the authentication data for TBRD may be compromised with trade-offs to system security. The SHA256 HMAC provided 256 bits of security, while the TESLA protocol used by Galileo operates with a truncated HMAC of only 40 bits (european_spce_agency_galileo_2021). If we compare the use of an equally secure digital signature scheme like ECC, the signature size would double the necessary length of the authentication data. When operating in a wireless environment with a high potential for packet loss, it is crucial to minimize the overhead required for message authentication. TBRD excels in this regard as the same message size can be authenticated with a smaller amount of data compared to a digital signature approach.
6. Discussion
@@
TBRD Parameter Configuration: The TBRD protocol supports several adjustable parameters to balance security and performance, including the delay disclosure period, interval period, and HMAC size. TESLA introduces a controlled delay $d$ between message transmission and key disclosure, which must be balanced against real-time needs. In latency-sensitive applications (e.g., drone command and control), longer delays degrade responsiveness, while shorter ones limit time for key distribution and verification. The interval duration $T_{int}$ should align with message frequency and network conditions: high-frequency messaging may require shorter $T_{int}$ to reduce overlap and buffer demand, while longer intervals reduce overhead but increase exposure to packet loss or manipulation. Optimal parameter tuning is crucial for striking a balance between security and operational efficiency. The HMAC output may also be truncated to accommodate lossy wireless channels.
@@
Mobile compromise: In TBRD, the key commitment of the TESLA-based broadcast authentication depends on the mobile device to serve as a trusted execution platform for key disclosure and signature operations. However, the integrity of this process hinges on the security guarantees of the TEE itself. Current TEEs (e.g., ARM TrustZone or Apple Secure Enclave) offer limited visibility and control to developers, and vendor-specific APIs or OS policies often constrain their use. Moreover, if the mobile device is compromised (e.g., rooted or jailbroken), the adversary may gain access to pre-disclosure keys. This highlights a significant limitation: the security model assumes the mobile TEE is trustworthy, but practical deployments must account for the possibility of device-level compromise.
@@
Standards integration: Integrating TESLA into the ASTM Remote ID framework presents both opportunities and challenges. TESLA offers a lightweight, broadcast-friendly authentication primitive that complements the ASTM requirement for message authenticity without necessitating public key infrastructure on resource-constrained drones. However, compliance with current ASTM standards (e.g., F3411) requires strict adherence to message construction and transmission. Furthermore, compatibility with various transmission mechanisms (Bluetooth, Wi-Fi) adds additional constraints.
7. Related Work
@@
Several works have focused on the privacy and security issues associated with compliance with the Remote ID rule. In  (tedeschi_privacy-aware_2024), Tedeschi et al. provide perhaps the most comprehensive overview of privacy and security issues facing Remote ID compliance by UAS operators and manufacturers. Their work identifies several potential threats and solutions, considering the security, privacy, and energy trade-offs associated with these solutions. While a wide variety of technical approaches were considered, their work did not address or mention the use of the TESLA algorithm.
@@
Of the numerous works that propose a technical solution or protocol to enhance Remote ID security, the most closely related is CAPSID (li_capsid_2024). In their work, Li et al. propose the CAPSID protocol, which builds upon the proposed use of session IDs. At its core, CAPSID utilizes certificates to authenticate broadcast messages. Our solution leverages the TESLA protocol to achieve authentication in a more computationally efficient manner. Additionally, they do not provide a practical solution to UAS capture. TBRD ensures that capture of a UAS does not give an adversary access to long-term key material. Furthermore, other closely related works attempt to provide solutions that enhance operator privacy. Privacy of operator location was not a focus of our work (li_capsid_2024; tedeschi_arid_2021; wisse_2_2023).
@@
The application of TESLA to solve current and future issues facing broadcast authentication is not unique. TESLA is in use by the Galileo global navigation satellite sytem (european_gnss_supervisory_authority_galileo_2022). TESLA was proposed as a solution for providing authentication to ADS-B in (berthier_sat_2017), which is widely used in the aviation industry. For the maritime sector, it has been proposed to enhance the authentication of Automated Identification Systems (AIS) (sciancalepore_auth-ais_2022) as well as its replacement, VHF Data Exchange (VDES)  (wimpenny_pragmatic_2025). Our system has many similarities with AuthAIS. One key difference is that AuthAIS utilizes a Trusted Third Party Maritime Authority to assign and distribute root keys to vessels that wish to use TESLA-authenticated AIS. They also propose a probabilistic security framework that reduces message overhead. Using TESLA to authenticate messages in these industries has not yet gained widespread adoption among the mainstream. The rapid expansion and constraints of UAS operations demand new solutions; our approach unites TESLA and trusted computing to surpass earlier limitations.
8. Conclusion